n! * (n+2)! = (2n)!
Find n=??

Answer by Edwin McCravy(20054)


It can't be a very large number since the right side
would be tremendously larger than the right side if n
were large.  So let's try n=1

1! * (1+2)! = (2∙1)!
1 * 3! = (2)!
1 * 6 = 2
6 = 2

No, that's false. So let's try n=2

2! * (2+2)! = (2∙2)!
2 * 4! = 4!
2 * 24 = 24
48 = 24

No, that's false. So let's try n=3

3! * (3+2)! = (2∙3)!
6 * 5! = 6!
6 * 120 = 720
720 = 720

Yes, that's true!  So n=3.

Edwin